If the average worker produces $55,000 of GDP, by how much will GDP change if there are 120 million labor force participants and the unemployment rate changes from 9.0 8.0 percent?


1
Expert's answer
2021-03-23T08:26:39-0400

number of unemployed persons at 9 percent"=" labor force"\times" unemployment rate"=120million\times0.09=10.8million"

At 8 percent, number unemployed"=120million\times0.08=9.6million"

change in unemployed persons"=10.8million-9.6million=1.2million=1,200,000"

Increase in GDP"=" production of average worker"\times" newly employed persons"=55,000\times 1,200,000=\$66 billion"